$SPAD/src/input palette.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- test for Palette domain
dark blue()
dim blue()
bright blue()
pastel blue()
light blue()
draw(x**2,x=-1..1,curveColor == hue dark blue())
shade bright blue()

———-

References^{[1] nothing}

4